OpenStudy (anonymous):

Which statement best explains whether y = 2x − 3 is a linear function or a nonlinear function?

OpenStudy (anonymous):

I have, 3 more of these. I just need a bit of help.

OpenStudy (kmeis002):

Are there statements or are you expected to write one?

OpenStudy (anonymous):

Which statement best explains whether y = 2x − 3 is a linear function or a nonlinear function? It is a linear function because the graph contains the points (0, −3), (1, −1), (2, 1), which are on a straight line. It is a nonlinear function because the graph contains the points (0, −3), (1, −1), (2, 1), which are not on a straight line. It is a linear function because the graph contains the points (−3, 0), (−1, 1), (1, 2), which are on a straight line. It is a nonlinear function because the graph contains the points (−3, 0), (−1, 1), (1, 2), which are not on a straight line.

OpenStudy (kmeis002):

It is indeed a linear function, but now you have to verify which sets of points fall on that line (hint, check them via the equation)

OpenStudy (anonymous):

Okay. That makes sence. I have 2 more. Which of the following has a graph that is a straight line? Equation 1: y = 4x3 + 6 Equation 2: y = 5x − 4.5 Equation 3: y2 = x − 1 Equation 4: y = 2x2 + 6 Thanks for your help again. :)

OpenStudy (kmeis002):

A line has the general form of \( y = mx+b\) (point slope).

OpenStudy (anonymous):

Okay, So I wasn't sure how to find if it was a straight line or not. :(